OpenStudy (anonymous):

An escalator in a department store is to carry people a vertical distance of 18 feet between floors. How long is the escalator if it makes and angle of 40° with the floor? Round to 1 decimal place. Select one: a. 9.6 ft b. 11.5 ft c. 17.9 ft d. none of the above

OpenStudy (anonymous):



OpenStudy (anonymous):

you are going to need a calculator for this \[\sin(40^\circ)=\frac{18}{x}\] so \[x=\frac{18}{\sin(40^\circ)}\]

OpenStudy (anonymous):

actually i never use a calculator i use this http://www.wolframalpha.com/input/?i=18%2Fsin%2840%29

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

x=28

OpenStudy (anonymous):

that is what i get yes

OpenStudy (anonymous):

then whats next

OpenStudy (anonymous):

idk have a beer?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

oh, i see since 28 is not a choice, pick D

OpenStudy (anonymous):

thats wrong that what i put

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i am sticking with \[\frac{18}{\sin(40^\circ)}\] and wolfram doesn't lie

OpenStudy (anonymous):

yeah i see still sticking to 28

OpenStudy (anonymous):

unless "how long is the escalator" means this 

OpenStudy (anonymous):

but that is not the length of the escalator that is the horizontal distance

OpenStudy (anonymous):

@jim_thompson5910 do you know?

jimthompson5910 (jim_thompson5910):

I'm getting 28.0030288834874 as well

jimthompson5910 (jim_thompson5910):

the horizontal component is roughly 21.4515646666956

jimthompson5910 (jim_thompson5910):

so this problem has a typo somewhere

OpenStudy (anonymous):

i guess there is i just wanted a second opinion

jimthompson5910 (jim_thompson5910):

I guess if there are no typos, and they intentionally put bad answers up there, then the answer would be D) none of the above. The others are probably put in there to throw you off?

OpenStudy (anonymous):

i put D though and got it wrong